# Quantum Fisher information matrices from Rényi relative entropies

### v1

@@ 9 Information matrices of classical–quantum states
In this section, I show how the various information matrices considered in this paper decompose whenever the unlderying parameterized family of states have a classical–quantum structure, generalizing the previous finding reported for the single-parameter case in [KW21, Proposition 7]. In particular, the various information matrices decompose as a sum of a classical part and a quantum part. For the purposes of this section, I adopt a revised notation in which the family on which the information matrix is being evaluated is explicitly written. That is, rather than the abbreviated notation used in (2.18), I write the following instead:
@@ 10 Conclusion
In summary, this paper explored quantum generalizations of the Fisher information matrix that are derived from quantum Rényi relative entropies, including the log-Euclidean, geometric, and $\alpha$-$z$ Rényi relative entropies. Among the various results of this paper, key contributions include a detailed proof of the formula in Theorem 10 for the $\alpha$-$z$ information matrix, the formula in Theorem for the $\alpha$-$z$ information matrix of parameterized thermal states, and orderings for the Petz– and sandwiched Rényi information matrices (Theorem 27 and Theorem 29).
Going forward from here, it is an interesting open question to determine the full range of $\alpha,z>0$ for which the data processing inequality holds for the $\alpha$-$z$ information matrix. Fact 9 provides sufficient conditions on $\alpha,z>0$ for which it holds, but Theorem 3 implies that these conditions are not necessary. By [Pet96, Theorems 3 and 5], determining the full range of $\alpha,z>0$ for which data processing holds is equivalent to determining the full range of $\alpha,z>0$ for which the function in Corollary 13 is operator monotone.
@@ Appendix B Review: Matrix derivatives from divided differences
This appendix provides a review of the method of divided differences for calculating matrix derivatives. Let $A(x)$ be a Hermitian matrix parameterized by $x\in\mathbb{R}$. Suppose that $A(x)$ has a spectral decomposition as follows:
@@ B.2 Examples
We now use Propositions 41 and 42 to find an integral formula for the derivative of all power functions.
@@
Applying Propositions 41 and 42, consider that
@@
Consider that
$\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{x-y}\right)\left(\frac{x^{\frac{\alpha}{z}}-y^{\frac{\alpha}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)$ $=\left(\frac{y^{\frac{1-\alpha}{z}}}{y}\right)\left(\frac{\left(\frac{x}{y}\right)^{\frac{1-\alpha}{z}}-1}{\frac{x}{y}-1}\right)\left(\frac{y^{\frac{\alpha}{z}}}{y^{\frac{1}{z}}}\right)\left(\frac{\left(\frac{x}{y}\right)^{\frac{\alpha}{z}}-1}{\left(\frac{x}{y}\right)^{\frac{1}{z}}-1}\right)$ (C.1)
$=\left(\frac{y^{\frac{1-\alpha}{z}}}{y}\right)\left(\frac{y^{\frac{\alpha}{z}}}{y^{\frac{1}{z}}}\right)\left(\frac{\left(\frac{x}{y}\right)^{\frac{1-\alpha}{z}}-1}{\frac{x}{y}-1}\right)\left(\frac{\left(\frac{x}{y}\right)^{\frac{\alpha}{z}}-1}{\left(\frac{x}{y}\right)^{\frac{1}{z}}-1}\right)$ (C.2)
$=\left(\frac{1}{y}\right)\left(\frac{\left(\frac{x}{y}\right)^{\frac{1-\alpha}{z}}-1}{\frac{x}{y}-1}\right)\left(\frac{\left(\frac{x}{y}\right)^{\frac{\alpha}{z}}-1}{\left(\frac{x}{y}\right)^{\frac{1}{z}}-1}\right).$ (C.3)
Now consider that
$\lim_{x\to y}\left(\frac{\left(\frac{x}{y}\right)^{\frac{1-\alpha}{z}}-1}{\frac{x}{y}-1}\right)\left(\frac{\left(\frac{x}{y}\right)^{\frac{\alpha}{z}}-1}{\left(\frac{x}{y}\right)^{\frac{1}{z}}-1}\right)$ $=\lim_{x\to 1}\left(\frac{x^{\frac{1-\alpha}{z}}-1}{x-1}\right)\left(\frac{x^{\frac{\alpha}{z}}-1}{x^{\frac{1}{z}}-1}\right)$ (C.4)
$=\left(\lim_{x\to 1}\frac{x^{\frac{1-\alpha}{z}}-1}{x-1}\right)\left(\lim_{x\to 1}\frac{x^{\frac{\alpha}{z}}-1}{x^{\frac{1}{z}}-1}\right)$ (C.5)
$=\left(\lim_{x\to 1}\frac{\frac{1-\alpha}{z}x^{\frac{1-\alpha}{z}-1}}{1}\right)\left(\lim_{x\to 1}\frac{\frac{\alpha}{z}x^{\frac{\alpha}{z}-1}}{\frac{1}{z}x^{\frac{1}{z}-1}}\right)$ (C.6)
$=\frac{\left(\frac{1-\alpha}{z}\right)\left(\frac{\alpha}{z}\right)}{\frac{1}{z}}\lim_{x\to 1}x^{\frac{1-\alpha}{z}-1+\frac{\alpha}{z}-1-\left(\frac{1}{z}-1\right)}$ (C.7)
$=\frac{\alpha\left(1-\alpha\right)}{z}\lim_{x\to 1}x^{-1}$ (C.8)
$=\frac{\alpha\left(1-\alpha\right)}{z}.$ (C.9)
Thus,
$\lim_{x\to y}\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{x-y}\right)\left(\frac{x^{\frac{\alpha}{z}}-y^{\frac{\alpha}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)=\alpha\left(\frac{1-\alpha}{z}\right)\left(\frac{1}{y}\right).$ (C.10)
Appendix D Proof of Equation (5.112)
@@
$=\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{\left(x-y\right)^{2}}\right)\left[\left(xy\right)^{\frac{\alpha}{z}}\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\left(x^{\frac{z-1+\alpha}{z}}-y^{\frac{z-1+\alpha}{z}}\right)\right]$ (D.1)
$=\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{x-y}\right)\left(\frac{1}{x-y}\right)\left[\left(xy\right)^{\frac{\alpha}{z}}\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\left(x^{\frac{z-1+\alpha}{z}}-y^{\frac{z-1+\alpha}{z}}\right)\right].$ (D.2)
@@
$=\left(\frac{\left(xy\right)^{\frac{\alpha}{z}}}{x-y}\right)\left[\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\left(x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)\right]$ (D.3)
$=\left(\frac{\left(xy\right)^{\frac{\alpha}{z}}}{x-y}\right)\left[\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\frac{\left(x^{\frac{1}{z}}-y^{\frac{1}{z}}\right)\left(x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right]$ (D.4)
@@
$\qquad\left[\left(x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\left(x^{\frac{1}{z}}-y^{\frac{1}{z}}\right)\left(x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)\right]$ (D.5)
Focusing on the term in the bottom line of (D.5), observe that
@@
$\qquad-\left(x^{\frac{1}{z}}\left(x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)-y^{\frac{1}{z}}\left(x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)\right)$ (D.6)
@@
$\qquad-\left(x^{\frac{1}{z}}x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-x^{\frac{1}{z}}y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}-y^{\frac{1}{z}}x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}+y^{\frac{1}{z}}y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)$ (D.7)
@@
$\qquad-\left(xy^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{\frac{1-\alpha}{z}}-x^{\frac{z-1}{z}}y^{\frac{1-\alpha}{z}}+yx^{-\frac{\alpha}{z}}\right)$ (D.8)
$=\left(xx^{-\frac{\alpha}{z}}+yy^{-\frac{\alpha}{z}}\right)-\left(xy^{-\frac{\alpha}{z}}+yx^{-\frac{\alpha}{z}}\right)$ (D.9)
$=\left(x-y\right)\left(x^{-\frac{\alpha}{z}}-y^{-\frac{\alpha}{z}}\right).$ (D.10)
Plugging back into (D.5), we conclude that
@@
$=\left(\frac{\left(xy\right)^{\frac{\alpha}{z}}}{\left(x-y\right)\left(x^{\frac{1}{z}}-y^{\frac{1}{z}}\right)}\right)\left(x-y\right)\left(x^{-\frac{\alpha}{z}}-y^{-\frac{\alpha}{z}}\right)$ (D.11)
$=\frac{\left(xy\right)^{\frac{\alpha}{z}}\left(x^{-\frac{\alpha}{z}}-y^{-\frac{\alpha}{z}}\right)}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}$ (D.12)
$=-\left(\frac{x^{\frac{\alpha}{z}}-y^{\frac{\alpha}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right).$ (D.13)
@@
$\left(\frac{1}{x-y}\right)\left[\left(xy\right)^{\frac{\alpha}{z}}\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\left(x^{\frac{z-1+\alpha}{z}}-y^{\frac{z-1+\alpha}{z}}\right)\right]\\ =-\left(\frac{x^{\frac{\alpha}{z}}-y^{\frac{\alpha}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right).$ (D.14)
Finally, plugging back into (D.2), we conclude that
$\left(xy\right)^{\frac{\alpha}{z}}\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{x-y}\right)^{2}-\left(\frac{x^{\frac{z-1+\alpha}{z}}-y^{\frac{z-1+\alpha}{z}}}{x-y}\right)\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{x-y}\right)\\ =-\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{x-y}\right)\left(\frac{x^{\frac{\alpha}{z}}-y^{\frac{\alpha}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right),$ (D.15)

### v2

@@ 10 Information matrices of classical–quantum states
In this section, I show how the various information matrices considered in this paper decompose whenever the underlying parameterized family of states have a classical–quantum structure, generalizing the previous finding reported for the single-parameter case in [KW21, Proposition 7]. In particular, the various information matrices decompose as a sum of a classical part and a quantum part. For the purposes of this section, I adopt a revised notation in which the family on which the information matrix is being evaluated is explicitly written. That is, rather than the abbreviated notation used in (2.18), I write the following instead:
@@ 11 Conclusion
In summary, this paper explored quantum generalizations of the Fisher information matrix that are derived from quantum Rényi relative entropies, including the log-Euclidean, geometric, and $\alpha$-$z$ Rényi relative entropies. Among the various results of this paper, key contributions include a detailed proof of the formula in Theorem 10 for the $\alpha$-$z$ information matrix, the formula in Theorem 32 for the $\alpha$-$z$ information matrix of parameterized thermal states, the formula in Theorem 35 for the $\alpha$-$z$ information matrix of time-evolved states, and orderings for the Petz– and sandwiched Rényi information matrices (Theorem 28 and Theorem 30).
Going forward from here, it is an interesting open question to determine the full range of $\alpha,z>0$ for which the data processing inequality holds for the $\alpha$-$z$ information matrix. Fact 9 provides sufficient conditions on $\alpha,z>0$ for which it holds, but Theorem 3 implies that these conditions are not necessary. By [Pet96, Theorems 3 and 5], determining the full range of $\alpha,z>0$ for which data processing holds is equivalent to determining the full range of $\alpha,z>0$ for which the function in Corollary 14 is operator monotone. It is also an intriguing open question to determine if there is an appealing expression for the $\alpha$-$z$ information matrices of parameterized thermal states and time-evolved states, like those in Theorem 32 and Theorem 35, for $\alpha>1$.
@@ Appendix B Review: Matrix derivatives from divided differences
This appendix provides a review of the method of divided differences for calculating matrix derivatives. See also [Bha97, Section V.3] and [Hia10, Sections 2.2 and 2.3].
For a general $d$-dimensional matrix $A(x)$, the matrix derivative $\frac{\partial}{\partial x}A(x)$ is defined as
$\frac{\partial}{\partial x}A(x)\coloneqq\begin{bmatrix}\frac{\partial}{\partial x}a_{1,1}(x)&\frac{\partial}{\partial x}a_{1,2}(x)&\cdots&\frac{\partial}{\partial x}a_{1,d}(x)\\ \frac{\partial}{\partial x}a_{2,1}(x)&\frac{\partial}{\partial x}a_{2,2}(x)&&\frac{\partial}{\partial x}a_{2,d}(x)\\ \vdots&&\ddots&\vdots\\ \frac{\partial}{\partial x}a_{d,1}(x)&\frac{\partial}{\partial x}a_{d,2}(x)&\cdots&\frac{\partial}{\partial x}a_{d,d}(x)\end{bmatrix}$. (B.1)
where the matrix elements of $A(x)$ are denoted by $a_{i,j}(x)$. From this, one can deduce that the product rule holds for matrices $A(x)$ and $B(x)$:
$\frac{\partial}{\partial x}\left[A(x)B(x)\right]=\left[\frac{\partial}{\partial x}A(x)\right]B(x)+A(x)\left[\frac{\partial}{\partial x}B(x)\right].$ (B.2)
B.1 First derivative
Let $A(x)$ be a Hermitian matrix parameterized by $x\in\mathbb{R}$. Suppose that $A(x)$ has a spectral decomposition as follows:
@@ B.2 Examples
We now use Propositions 47 and 48 to find an integral formula for the derivative of all power functions.
@@
Applying Propositions 47 and 48, consider that
@@
$=\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{\left(x-y\right)^{2}}\right)\left[\left(xy\right)^{\frac{\alpha}{z}}\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\left(x^{\frac{z-1+\alpha}{z}}-y^{\frac{z-1+\alpha}{z}}\right)\right]$ (C.1)
$=\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{x-y}\right)\left(\frac{1}{x-y}\right)\left[\left(xy\right)^{\frac{\alpha}{z}}\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\left(x^{\frac{z-1+\alpha}{z}}-y^{\frac{z-1+\alpha}{z}}\right)\right].$ (C.2)
@@
$=\left(\frac{\left(xy\right)^{\frac{\alpha}{z}}}{x-y}\right)\left[\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\left(x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)\right]$ (C.3)
$=\left(\frac{\left(xy\right)^{\frac{\alpha}{z}}}{x-y}\right)\left[\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\frac{\left(x^{\frac{1}{z}}-y^{\frac{1}{z}}\right)\left(x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right]$ (C.4)
@@
$\qquad\left[\left(x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\left(x^{\frac{1}{z}}-y^{\frac{1}{z}}\right)\left(x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)\right]$ (C.5)
Focusing on the term in the bottom line of (C.5), observe that
@@
$\qquad-\left(x^{\frac{1}{z}}\left(x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)-y^{\frac{1}{z}}\left(x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)\right)$ (C.6)
@@
$\qquad-\left(x^{\frac{1}{z}}x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}-x^{\frac{1}{z}}y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}-y^{\frac{1}{z}}x^{\frac{z-1}{z}}y^{-\frac{\alpha}{z}}+y^{\frac{1}{z}}y^{\frac{z-1}{z}}x^{-\frac{\alpha}{z}}\right)$ (C.7)
@@
$\qquad-\left(xy^{-\frac{\alpha}{z}}-y^{\frac{z-1}{z}}x^{\frac{1-\alpha}{z}}-x^{\frac{z-1}{z}}y^{\frac{1-\alpha}{z}}+yx^{-\frac{\alpha}{z}}\right)$ (C.8)
$=\left(xx^{-\frac{\alpha}{z}}+yy^{-\frac{\alpha}{z}}\right)-\left(xy^{-\frac{\alpha}{z}}+yx^{-\frac{\alpha}{z}}\right)$ (C.9)
$=\left(x-y\right)\left(x^{-\frac{\alpha}{z}}-y^{-\frac{\alpha}{z}}\right).$ (C.10)
Plugging back into (C.5), we conclude that
@@
$=\left(\frac{\left(xy\right)^{\frac{\alpha}{z}}}{\left(x-y\right)\left(x^{\frac{1}{z}}-y^{\frac{1}{z}}\right)}\right)\left(x-y\right)\left(x^{-\frac{\alpha}{z}}-y^{-\frac{\alpha}{z}}\right)$ (C.11)
$=\frac{\left(xy\right)^{\frac{\alpha}{z}}\left(x^{-\frac{\alpha}{z}}-y^{-\frac{\alpha}{z}}\right)}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}$ (C.12)
$=-\left(\frac{x^{\frac{\alpha}{z}}-y^{\frac{\alpha}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right).$ (C.13)
@@
$\left(\frac{1}{x-y}\right)\left[\left(xy\right)^{\frac{\alpha}{z}}\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}\right)-\left(x^{\frac{z-1+\alpha}{z}}-y^{\frac{z-1+\alpha}{z}}\right)\right]\\ =-\left(\frac{x^{\frac{\alpha}{z}}-y^{\frac{\alpha}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right).$ (C.14)
Finally, plugging back into (C.2), we conclude that
$\left(xy\right)^{\frac{\alpha}{z}}\left(\frac{x^{\frac{z-1}{z}}-y^{\frac{z-1}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right)\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{x-y}\right)^{2}-\left(\frac{x^{\frac{z-1+\alpha}{z}}-y^{\frac{z-1+\alpha}{z}}}{x-y}\right)\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{x-y}\right)\\ =-\left(\frac{x^{\frac{1-\alpha}{z}}-y^{\frac{1-\alpha}{z}}}{x-y}\right)\left(\frac{x^{\frac{\alpha}{z}}-y^{\frac{\alpha}{z}}}{x^{\frac{1}{z}}-y^{\frac{1}{z}}}\right),$ (C.15)